OpenStudy (mathmath333):

In how many ways can 945 be written as difference of squares of two natural numbers ?

OpenStudy (anonymous):

I got one way : \(31^2 - 4^2\) :P

mathslover (mathslover):

5 = 6 - 1 5 = 7 - 2 5 = 8 - 3 5 = 9 - 4 So, two cases possible.. 5 = 6 - 1 or 5 = 9- 4 \(945 = (2k+1)^2 - (2k)^2 \) \(945 = (2k+1-2k)(2k+1+2k)\) \(945=(4k+1)\) \(4k = 944\) \(k = 236\) So, one case is this when k = 236 i.e the natural numbers are 472 and 473

ganeshie8 (ganeshie8):

i found 8 ways

mathslover (mathslover):

Now, let us consider that the two natural numbers are not consecutive. \(945 = (2m + 1)^2 - (2n)^2 \) \(945 = 4m^2 + 1 + 4m - 4n^2\) \(945 = (4m^2 - 4n^2) + 4m + 1\) \(944 = 4(m^2 - n^2 + m) \\ 236 = m^2 - n^2 + m \)

OpenStudy (mathmath333):

yes thats right , but how ?

OpenStudy (mathmath333):

i m only concerned with the number of ways , the exact values are not relevant

mathslover (mathslover):

Oh, I see. Then there must be a specific algorithm for it. Would love to see how ganeshie approached the solution.

ganeshie8 (ganeshie8):

Look at prime factorization \[945 = 3^3\times 5\times 7 = x^2-y^2 = (x+y)(x-y)\]

mathslover (mathslover):

No, don't tell me that! Ganeshie... you just wasted a tree... here is how i) I saw the problem. ii) I started working on it. iii) There was no one else posting answer so I resumed my work. iv) I finished a page of work. v) Then Ganeshie posted the solution. vi) That one page got wasted. vii) And in this way, one tree got consumed! :'( Jokes apart, smartly done @ganeshie8

ganeshie8 (ganeshie8):

Its all good :) Own solution is always better than somebody else's solution :) it tells us alternative ways to look at the problem sometimes... but what i gave was just a hint...

ganeshie8 (ganeshie8):

\[\large 3^{\color{Red}{3}}\times 5^{\color{red}{1}}\times 7^{\color{red}{1}} = (x+y)(x-y) \] Next notice that \(945\) has a total of \((\color{Red}{3}+1)(\color{Red}{1}+1)(\color{Red}{1}+1) = 16\) divosors. But we are only interested in positive solutions : \(x\gt 0 \) and \(y \gt 0\)

OpenStudy (mathmath333):

perfect!

ganeshie8 (ganeshie8):

Suppose \(x+y = d\) then we have \(x-y = \dfrac{945}{d}\) solving gives \[x = \dfrac{d+\frac{945}{d}}{2}\] \[y = \dfrac{d-\frac{945}{d}}{2}\]

ganeshie8 (ganeshie8):

Notice that \(x\) is positive for all divosors \(d\), but \(y\) is positive only for half of the divisors. Consequently the number of solutions will be half of the total number of positive divisors of \(945\)

OpenStudy (mathmath333):

thnks got it